OpenStudy (anonymous):

10a^2 25a^3 ------- +------- 7b^2 14b^3 it is asking to be simplify. and I don't know how to do that:)

OpenStudy (anonymous):

\[\frac{ 10a^2 }{ 7b^2 }+\frac{ 25a^3 }{ 14b^3 }\] multiply tha numerator and denominator of \[\frac{ 10a^2 }{ 7b^2 }\] by \[2b\] that will make you have a common denominator

OpenStudy (anonymous):

then i think you can do the rest yourself